You cross a Drosophila female with kidney-shaped and brown eyes with a male that has wildtype eye shape and color. Kidney-shaped eyes are found in females that are heterozygous for the bar eyed mutation. Bar eyed females are homozygous for the bar eyed mutant allele, whereas bar eyed males are hemizygous for the mutation. Brown eyes are an autosomal recessive mutation. Attention: Half of the resulting F1 has brown eyes!

                           [ Select ]                       ["1/4", "1/2", "1/12", "1/8", "1/16", "1/24", "1/3", "0"]           of the F1 will be females with wildtype eye shape and color
Answer #1

The answer will be 1/16.

Explanation: Let, B = allele for wild type color, b = allele for brown color

As brown eyes due to autosomal recessive mutation & cross between brown eyed female & wild type eyed male produces half of the F1 as brown eyed, wild type male is heterozygous for wild type color, i.e., genotype Bb. Brown eyed female will have bb genotype. The cross between them is shown in below Punnett square.

Gametes b b
B Bb (Wild type color) Bb (Wild type color)
b bb (Brown color) bb (Brown color)

Now, as possibility of male & female offspring is equal, i.e., each has 1/2 probability of occurring, we will expect to observe wild type & brown color in F1 male & female in equal ratio. So, from the above Punnett square, we find that probability of F1 female with wild type eye color is 1/4 (As 1/2 of F1 will wild type color & from that half will be female).

According to the question, eye shape is sex linked.

Let, X = wild type X chromosome, X' = X chromosome responsible for bar eyed mutation

Now, XX' genotype will have kidney shaped eye in female, X'X' will have bar eyed eye in female & X'Y will have bar-eyed eye in male.

As Drosophila female has kidney-shaped eye, her genotype will be XX'. Male will have XY genotype. The cross between them is shown in below Punnett square.

Gametes X X'
X XX (Wild type female) XX' (Kidney shaped female)
Y XY (Wild type male) X'Y (Bar-eyed male)

From the above Punnett square, we find that probability of F1 female with wild type eye shape is 1/4.

So, probability of F1 females with wild type eye shape & color = Probability of F1 female with wild type eye color x Probability of F1 female with wild type eye shape = 1/4 x 1/4 = 1/16